# Density-aware Global-Local Attention Network for Point Cloud Segmentation

### v1

@@ Abstract
I Introduction
@@
Figure 1: Global attention carried out between local areas and local attention with different window sizes within local areas of different densities in proposed method.
@@
The main contributions of this paper are:
  • A novel density-aware local attention module is proposed. In this module, we determine the window size based on the density distribution, the higher the density, the smaller the window and vice versa.
  • We present a global attention module based on performing inter-area feature updates. It is designed to model long-range dependencies where each local area is considered as an independent whole.
  • We introduce the category response loss. We use the category information embedded in the scene to constrain the network in the middle layer so that the network can respond to the presence of categories in advance.
  • Our method obtains excellent experimental results on both public datasets and self-collected datasets of real scenes. In experiments on the commonly used indoor dataset S3DIS [8], our method achieves excellent segmentation results in most categories.
II Related Work
II-A 3D Point Cloud Segmentation
The goal of the 3D point cloud segmentation task is to process the point cloud input to the network to obtain the category labels corresponding to each 3D point. Commonly used 3D point cloud segmentation methods can be classified into the following categories based on the principles of the network model and the data format of the input network: voxel-based methods [9, 10], projection-based methods [11, 12, 13, 14], MLP-based methods [15, 16, 17, 18, 19], pointwise convolution-based methods [20, 21], methods combining image processing [22, 23, 24], and methods based on the attention mechanism [2, 5, 25, 6, 26, 27, 28, 29, 30, 31, 7, 3, 4, 32, 33]. Voxel-based approaches discretize the 3D space into regular grids to enable 3D convolutions, but often struggle with trade-offs between resolution and efficiency. Projection-based methods transform 3D points into 2D representations, leveraging mature 2D CNN architectures for feature learning. MLP-based methods, such as PointNet and its successors, operate directly on raw point clouds, introducing innovative mechanisms to handle unordered data. Recent advances in pointwise convolution and attention mechanisms have further enhanced the ability to capture local and global geometric relationships.
II-B Network based on Transformer
@@
Figure 3: Visualisation diagram of local area density division of point cloud data (the visualisation of point cloud file density division in the indoor dataset S3DIS [8].
Most of the current point cloud transformer networks [5, 25, 6, 26] only consider local features without considering long-range dependencies, and are unable to better extract complete global context information, and including the above attention methods based on the transformer network architecture, although it expands the range of sensory field and information interaction to a certain extent, part of the long-range information is still not considered, and the window division also affects the performance. In real scenes such as oil field production, there are a large number of empty or weakly textured areas, and at the same time, there are small valves, railings, pipelines and other slender objects, and it is difficult for existing methods to perform point cloud segmentation better. Therefore, the proposed method is designed in terms of local window segmentation strategy, global feature update path, and loss function to further solve the problems existing in the current network.
III Methodology
@@
III-A Density-based Global-Local Attention
@@
III-A1 Local Feature Updates
Conventional Transformer-based 3D point cloud semantic segmentation networks [2, 5, 25, 6, 26], due to the large computational resources they require, need to first divide the input point cloud into grids of uniform size and sample the panels when there are too many points in them for local feature updating. However, the density of the point cloud data varies in different local areas, and setting a uniform grid size will cause the dense area of the overall point cloud to discard more 3D points compared to the sparse area when sampling, resulting in the loss of too many 3D points of smaller objects in the dense area. In this regard, the overall point cloud is first divided into multiple local areas of the same size. Then we set the number of local areas to be divided, and calculate the size of the local areas according to the number of areas and the real space occupied by the input point cloud. After that, we count the number of 3D points in each local area, and calculate the density of each local area according to the volume of the corresponding area. After calculating the density of each local area, all areas are sorted from sparse to dense density, divided into multiple discrete parts of density. We then assign a density value to the corresponding local area and all 3D points contained within it, with the sparsest area having the smallest density value and vice versa.
We set the same window size for local areas belonging to the same discrete part of the point cloud density, setting larger windows for areas with smaller density values and vice versa. As shown in Fig. 3 is the visualisation of the indoor scene point cloud data processed using the proposed local area density division module. The different colours represent the different density degree of the corresponding local areas of the point cloud, and it can be seen that the local areas containing special categories (e.g., columns, cabinets, etc., which are boxed by purple) are distinctly different from the density degree assigned to the neighbouring areas.
@@ III-A1 Local Feature Updates
where $Linear$ represents the linear layer operation and its superscripts represent the input and output dimensions of this linear layer. The $attn_{i,j,h}$ is the attention value computed by dot product of the vectors corresponding to the $i$th Query point and the $j$th Key point of the $h$th attention head in the window. The $out$ is then the aggregated features obtained by the computation. Finally, the features are projected back to the dimension consistent with the input features through an additional linear layer to obtain the final output feature $o$.
@@
III-A2 Global Feature Updates
For the global feature update process, we regard each local area as a whole that aggregates the position and feature information of the corresponding area to calculate the attention map between different local areas. Before calculating the global attention, the information for local area needs to be aggregated, and this aggregation process is mainly divided into two steps, the first step is the aggregation of 3D spatial location information, as shown in Equation 5; the second step is the aggregation of feature information of the local area, as shown in Equation 6:
@@ III-B Loss Function
TABLE I: Quantitative comparison experiment and ablation experiment results of semantic segmentation of indoor point cloud dataset S3DIS [8] Area 5.
@@
where $C$ represents the total number of categories. $N$ represents the total number of points. $w_{k}$ represents the adaptive weights of the $k$th category in calculating the cross-entropy loss, which is inversely proportional to the ratio of the number of samples of the corresponding category points to the total number of points extracted from the partially randomly selected scenarios multiple times. $l_{i}k$ represents the labelling of whether the $i-$th point belongs to the $k-$th category or not, when the $i-$th point belongs to the $k-$th category, the corresponding $l_{ik}$ value is $1$, otherwise it is 00. The $p_{ik}$ represents the probability that the network predicts the $k-$th category for the $i-$th point.
@@
where $s_{ik}^{l}$ represents the binary classification labels of the $i-$th point against the $k-$th category, if the $i-$th point belongs to the $k-$th category, the corresponding $s_{ik}^{l}$ value is $1$, otherwise it is 00. The $s_{ik}^{p}$ represents the binary prediction result for the $i-$th point against the $k-$th category.
@@ IV Experiments
In order to verify the effectiveness of the proposed network approach, we choose to quantitatively compare it with existing representative methods on commonly used indoor 3D point cloud datasets, S3DIS [8], ScanNet [34], and qualitatively compares the segmentation results with the baseline network processing on an outdoor 3D point cloud dataset, Semantic3D [35]. The proposed network is also applied to the part segmentation task by using the ShapeNet [36] dataset for the corresponding experiments. In addition, we collect point cloud data from oil field production scenarios and perform segmentation experiments on these outdoor real field point clouds to verify the generality of the proposed method. The designed ablation experiments evaluate the effectiveness of the DbGLA module with the proposed CR-Loss.
IV-A Experimental settings
IV-A1 Comparison Methods
@@
In this paper, the proposed methods are compared with existing representative methods applied to the task of semantic segmentation of 3D point clouds, respectively. Specifically, we choose MLP-based methods (PointNet [15], PointNet++ [16], RandLA-Net [18]), point convolution methods (KPConv [21]), and other Transformer network-based methods for attention mechanisms, which include the latest state-of-the-art methods that have been published in recent years: Point Cloud Transformer [2], Point Transformer [5], Stratified Transformer [6], Superpoint Transformer [27], SPoTr [28], SuperCluster [29], Point Transformer V2 [26] and OctFormer [37].
IV-A2 Datasets
@@
In addition to the public dataset, we also conduct experiments in a real-scene dataset, which is the self-acquisition point cloud data collected from the real oil field scene by the air-ground fusion acquisition equipment consisting of 16-line radar, camera, UAV, IMU, etc. The point cloud involves a real space covering about 30,000 square metres, and contains more open weakly textured areas in the point cloud scene, as well as fewer samples of elongated objects in the categories of pipelines, fences, poles, road surfaces, containers, scaffolding, pumping machines, etc.
IV-A3 Experimental Details
Following the practice of new point cloud segmentation models in recent years [27, 28, 29], we use Overall Accuracy (OA), mean Accuracy (mAcc), mean Intersection over Union (mIoU) for semantic segmentation and Class mIoU, Instance mIoU for part segmentation. The deployment platforms for the experiments are all PyTorch, the computing platform for the indoor scene experiments is a server equipped with two Titan RTX (24G*2 video memory), and the computing platform for the outdoor scene experiments is a host computer equipped with an RTX4090 (24G video memory). The optimiser in the training process is selected as Adam, and the initial learning rate is set to 0.006. For the hyperparameter settings involved in the article, the number of discrete parts of the densities is generally set to 5 by default in the indoor scenario, while the number of discrete parts is set to 7 in the outdoor scenario; and the hyperparameter used in the loss function to balance the two loss terms is set to 0.5 by default. In addition, the number of iterations of the proposed method is set to 100 by default for experimental training.
IV-B Indoor Scene Experiment
@@
TABLE II: Quantitative comparison experiment results of semantic segmentation of indoor point cloud dataset S3DIS [8] (6-fold validation).
@@
Table I summarises the results for the indoor scene dataset S3DIS [8] Area 5 experimental scenarios. The experimental results are listed as overall and category-level accuracy metrics for the comparison methods [21, 15, 2, 6, 27, 28] and for the proposed method. Bolded data represent the best results for the corresponding metrics, while underlining represents the second best results for the metrics. In the overall indexes, they are better than other comparative methods and state-of-the-art methods, and the proposed method category intersection ratio indexes are also the best in six of the 13 categories, while the other seven categories have also reached the sub-optimal level. Table II summarises the results under the six-fold cross-validation experimental scenario for the indoor scene dataset S3DIS [8]. The overall accuracy, average accuracy, and average cross-validation ratio values of the comparison method [21, 15, 16, 18, 5, 29], as well as the method of this paper, are listed in the experimental results. In addition, Table III shows the experimental results on the validation set of the ScanNet [34] dataset.
TABLE III: Quantitative comparison experiment results of semantic segmentation of indoor point cloud dataset ScanNet [34].
@@
IV-C Outdoor Scene Experiment

### v2

@@ Abstract
keywords:
density awareness , point cloud segmentation , three-dimensional visual transformer network , global attention , category response loss
journal: Image and Vision Computing
1 Introduction
@@
Figure 1: Global attention carried out between local areas and local attention with different window sizes within local areas of different densities.
@@
Following the introduction, Section 2 (Related Work) reviews existing literature on 3D point cloud segmentation and Transformer-based networks. Section 3 (Methodology) details the proposed approach, introducing the Density-based Global-Local Attention mechanism (with subsections on local and global feature updates) and the novel loss function. Section 4 (Experiments) comprehensively evaluates the method, describing the experimental settings, and presenting results on indoor scene segmentation, outdoor scene segmentation, part segmentation, and a series of ablation studies. Finally, the paper concludes with Section 5 (Conclusions), summarizing the findings, contributions, limitation and future work.
The main contributions of this paper are:
  • 1.
    Our work pioneers the integration of local, density-aware attention and global attention. It dynamically reassigns feature aggregation boundaries via DBSCAN-based density partitioning, replacing static geometric partitioning.
  • 2.
    The proposed attention mechanism divides point clouds into discrete density parts, each of which is assigned a local attention window. This achieves higher accuracy by focusing computational resources intensively on detailed information in dense regions.
  • 3.
    The introduction of category-response loss (CR-loss) establishes an intermediate supervisory mechanism that enforces category-specific feature separation at the early stages of the network, collectively forming a distinct point cloud segmentation framework architecturally.
  • 4.
    Exhaustive comparison experiments on commonly used public datasets as well as self-acquisition datasets and ablation experiments show that our proposed method has excellent segmentation results, validating the effectiveness and robustness of our work.
2 Related Work
2.1 3D Point Cloud Segmentation
The goal of the 3D point cloud segmentation is to process the point cloud input to obtain the category labels corresponding to each 3D point. Commonly used 3D point cloud segmentation methods can be classified into the following categories based on the principles of the network and the data format of the input: voxel-based methods [8, 9], projection-based methods [10, 11, 12], MLP-based methods [13, 14, 15, 16, 17], pointwise convolution-based methods [18, 19, 20, 21], graph convolution-based methods [22, 23], methods combining image processing [24, 25], and attention-based methods [2, 3, 4, 5, 6, 7, 26, 27, 28, 29, 30, 20, 31].
SSCNs [8] introduces Submanifold Sparse Convolutional Networks, which efficiently process sparse 3D point clouds by restricting convolutions to non-empty voxels for semantic segmentation. MinkowskiNet [9] proposes Minkowski Convolutional Neural Networks using 4D spatio-temporal sparse tensors to achieve large-scale spatiotemporal point cloud segmentation. These voxel-based methods discretise the 3D space into regular grids to enable 3D convolutions, but often struggle with trade-offs between resolution and efficiency. F. J. Lawin et al. [10] design a deep projective 3D segmentation framework by projecting point clouds to 2D multi-views and fusing features to enhance semantic segmentation accuracy. Sueezeseg [11] combines CNNs with CRF to project 3D LiDAR point clouds onto 2D spherical grids for real-time road-object segmentation. PolarNet [12] improves polar grid representation using annular partitioning to enhance LiDAR point cloud segmentation in bird’s-eye view. These projection-based methods transform 3D points into 2D representations, leveraging mature 2D CNN architectures for feature learning. MLP-based methods, such as PointNet [13] and its successors PointNet++ [14], DeepGCNs [15], RandLA-Net [16], ThickSeg [17], operate directly on raw point clouds, introducing innovative mechanisms to handle unordered data. Recent advances in point-wise convolution and attention mechanisms have further enhanced the ability to capture local and global geometric relationships.
@@ 2 Related Work
Figure 3: Visualisation diagram of local area density division of point cloud (the visualisation of point cloud file density division in SemanticKITTI [32].
2.2 Network based on Transformer
@@
Most of the current point cloud transformer networks [5, 6, 26, 27] only consider local features without considering long-range dependencies, and are unable to better extract complete global context information. Point Transformer [5] innovatively applies transformer-based self-attention mechanisms to 3D point clouds, adaptively modeling contextual relationships between points with geometrically weighted attention to capture local structures and enhance feature representation. Stratified Transformer [6] proposes a hierarchical architecture that processes 3D point clouds at multiple scales with stratified self-attention, capturing both fine-grained details and global contextual information for improved semantic segmentation. Fast Point Transformer [26] introduces an efficient framework that accelerates self-attention computation for point clouds via a hierarchical grouping strategy and approximate attention, enabling real-time processing while maintaining segmentation accuracy. Point Transformer V2 [27] advances the framework with grouped vector attention, which models point features as vector groups to capture directional relationships, and partition-based pooling, which enhances spatial context aggregation for more discriminative 3D feature learning. The above attention methods based on the transformer network architecture, although it expands the range of sensory field and information interaction to a certain extent, part of the long-range information is still not considered, and the window division also affects the performance. In real scenes such as oil field production, there are a large number of empty or weakly textured areas, and at the same time, there are small valves, railings, pipelines and other slender objects, and it is difficult for existing methods to perform point cloud segmentation better. Therefore, the proposed method is designed in terms of local window segmentation strategy, global feature update path, and loss function to further solve the problems existing in the current network.
3 Methodology
@@
3.1 Density-based Global-Local Attention
@@
3.1.1 Local Feature Updates
Conventional Transformer-based 3D point cloud semantic segmentation networks [2, 5, 6, 26, 27], due to the large computational resources they require, need to first divide the input point cloud into grids of uniform size and sample the panels when there are too many points in them for local feature updating. However, the density of the point cloud data varies in different local areas, and setting a uniform grid size will cause the dense area of the overall point cloud to discard more 3D points compared to the sparse area when sampling, resulting in the loss of too many 3D points of smaller objects in the dense area. In this regard, the overall point cloud is first divided into multiple local areas of the same size. Then we set the number of local areas to be divided, and calculate the size of the local areas according to the number of areas and the real space occupied by the input point cloud.
Then we use the idea of DBSCAN[33] to partition different local regions of the input point cloud according to different densities. The DBSCAN algorithm, a density-based clustering method, operates by first selecting an unvisited point in the dataset and identifying its neighboring points within a specified radius, $\epsilon$. If the number of neighboring points is greater than or equal to a predefined threshold, the point is marked as a core point and a new cluster is initiated. Then, the algorithm recursively expands the cluster by adding all density-reachable points to the cluster, which are those within the $\epsilon$ radius of the core point. This process continues, incorporating neighboring points and forming a dense cluster. If a point does not meet the density criteria, it is labeled as noise. The algorithm iterates through the dataset, repeating this process for each unvisited core point until all points have been visited and assigned to a cluster or labeled as noise. This approach allows DBSCAN to identify clusters of arbitrary shapes and effectively handle outliers. After partition different local areas of the input point cloud according to different densities, all areas are sorted from sparse to dense density, divided into multiple discrete parts of density. We then assign a density value to the corresponding local area and all 3D points contained within it, with the sparsest area having the smallest density value and vice versa.
We set the same window size for local areas belonging to the same discrete part of the point cloud density, setting larger windows for areas with smaller density values and vice versa. As shown in Fig. 3 is the visualisation of the outdoor scene point cloud data processed using the proposed local area density division module. The different colours represent the different density degree of the corresponding local areas of the point cloud, and it can be seen that the local areas containing special categories (e.g., car, motorcycle, etc., which are boxed by red) are distinctly different from the density degree assigned to the neighbouring areas.
@@ 3.1.1 Local Feature Updates
where $Linear$ represents the linear layer operation and its superscripts represent the input and output dimensions of this linear layer. The $attn_{i,j,h}$ is the attention value computed by dot product of the vectors corresponding to the $i$th Query point and the $j$th Key point of the $h$th attention head in the window. The outout is then the aggregated features obtained by the computation. Finally, the features are projected back to the dimension consistent with the input features through an additional linear layer to obtain the final output feature $o$.
@@
3.1.2 Global Feature Updates
For the global feature updates, we regard each local area as a whole that aggregates the position and feature information of the corresponding area to calculate the attention map between different local areas. Before calculating the global attention, the information for local area needs to be aggregated, and this aggregation process is mainly divided into two steps, the first step is the aggregation of 3D spatial location information, as shown in Equation 5; the second step is the aggregation of feature information of the local area, as shown in Equation 6:
@@ 3.2 Loss Function
Table 1: Quantitative comparison experiment and ablation experiment results of semantic segmentation of indoor point cloud dataset S3DIS [34] Area 5.
@@
where $C$ represents the total number of categories. $N$ represents the total number of points. $w_{k}$ represents the adaptive weights of the $k$th category in calculating the cross-entropy loss, which is inversely proportional to the ratio of the number of samples of the corresponding category points to the total number of points extracted from the partially randomly selected scenarios multiple times. $l_{ik}$ represents the labelling of whether the $i$-th point belongs to the $k$-th category or not, when the $i$-th point belongs to the $k$-th category, the corresponding $l_{ik}$ value is $1$, otherwise it is 0. The $p_{ik}$ represents the probability that the network predicts the $k$-th category for the $i$-th point.
@@
where $s_{ik}^{l}$ represents the binary classification labels of the $i$-th point against the $k$-th category, if the $i$-th point belongs to the $k$-th category, the corresponding $s_{ik}^{l}$ value is $1$, otherwise it is 0. The $s_{ik}^{p}$ represents the binary prediction result for the $i$-th point against the $k$-th category.
@@ 4 Experiments
In order to verify the effectiveness of the proposed method, we choose compare it with existing representative methods on several commonly used indoor and outdoor 3D point cloud datasets. The proposed network is also applied to the part segmentation task for the corresponding experiments. In addition, we collect point cloud data from oil field production scenarios and perform segmentation experiments on these outdoor real field point clouds to verify the generality of the proposed method. The designed ablation experiments evaluate the effectiveness of the DbGLA module with the proposed CR-Loss.
4.1 Experimental settings
4.1.1 Comparison Methods
@@
In this paper, the proposed methods are compared with existing representative methods applied to the task of semantic segmentation of 3D point clouds, respectively. Specifically, we choose MLP-based methods (PointNet [13], PointNet++ [14], RandLA-Net [16]), point convolution methods (KPConv [19], OA-CNNs [35], MinkUNet [36], SPVNAS [37], Cylinder3D [38]), methods combining image processing (WaffleIron [39], 2DPASS [40], FRNet [41]) and other Transformer network-based methods for attention mechanisms, which include the latest state-of-the-art methods that have been published in recent years: Point Cloud Transformer [2], Point Transformer [5], Stratified Transformer [6], Superpoint Transformer [28], SPoTr [29], SuperCluster [30], Point Transformer V2 [27], OctFormer [42], SphereFormer [43], Point Transformer V3 [44].
4.1.2 Datasets
Following the practice of new point cloud segmentation models in recent years, we conducted our semantic segmentation experiments on six popular benchmarks, including the indoor scene datasets S3DIS [34], ScanNet [45], as well as the outdoor scene datasets Semantic3D [46], SemanticKITTI [32], nuScenes [47] and Waymo Open [48]. S3DIS [34] contains 271 indoor rooms across 6 large-scale areas, providing point cloud data with 13 semantic class annotations (e.g., furniture, walls), tailored for 3D semantic and instance segmentation in indoor scenes. ScanNet [45] includes over 1500 indoor scans with 2.5 million RGB-D views, offering 3D camera poses, surface reconstructions, and instance-level semantic labels, supporting point cloud segmentation and scene understanding tasks. Semantic3D [46] comprises over 3 billion outdoor points across urban streets, churches, etc., labeled with 8 semantic classes (e.g., buildings, vegetation), designed for large-scale 3D semantic segmentation in complex outdoor environments. SemanticKITTI [32] extends KITTI with 22 sequences of LiDAR point clouds, distinguishing dynamic/static objects via dense annotations for 22 classes, enabling temporal point cloud segmentation in autonomous driving scenarios. NuScenes [47] captures multi-modal data (LiDAR, camera, radar) across 2 cities (Boston, Singapore), with its LiDAR Seg subset providing 32-class annotations for 23 object categories, supporting point cloud segmentation in complex traffic scenes. Waymo Open Dataset [48] integrates LiDAR, camera, and radar data across 5000+ driving segments (including 2025 additions), focusing on long-tail outdoor scenarios (e.g., construction zones), with semantic labels for point cloud segmentation in autonomous driving.
@@ 4.1.2 Datasets
In addition to the public dataset, we also conduct experiments in a real-scene dataset, which is the self-acquisition point cloud data collected from the real oil field scene by the air-ground fusion acquisition equipment consisting of 16-line radar, camera, UAV, IMU, etc. The point cloud involves a real space covering about 30,000 square metres, and contains more open weakly textured areas in the point cloud scene, as well as fewer samples of elongated objects in the categories of pipelines, fences, poles, road surfaces, containers, scaffolding, pumping machines, etc. In addition, we experimentally test the performance of the proposed method on part segmentation tasks on the ShapeNet [49] dataset. ShapeNet-Part [49] focuses on part segmentation with 16,881 models from 16 categories annotated with 50 part classes (e.g., airplane wings, chair legs), making it a benchmark for 3D point cloud part segmentation tasks.
@@
Table 2: Quantitative comparison experiment results of semantic segmentation of indoor point cloud dataset S3DIS [34] (6-fold validation).
@@
4.1.3 Experimental Details
Following the practice of new point cloud segmentation models in recent years [28, 29, 30], we use Overall Accuracy (OA), mean Accuracy (mAcc), mean Intersection over Union (mIoU) for semantic segmentation and Class (Cls.) mIoU, Instance (Ins.) mIoU for part segmentation. The deployment platforms for the experiments are all PyTorch, the computing platform for the indoor scene experiments is a server equipped with two Titan RTX (24G*2 video memory), and the computing platform for the outdoor scene experiments is a host computer equipped with an RTX4090 (24G video memory). The optimiser in the training process is selected as Adam, and the initial learning rate is set to 0.006. For the hyperparameter settings involved in the article, the number of discrete parts of the densities is generally set to 5 by default in the indoor scenario, while the number of discrete parts is set to 7 in the outdoor scenario. And the hyperparameter used in the loss function to balance the two loss terms is set to 0.5 by default, in order to prevent either loss component from dominating, while optimally exploiting their complementary strengths in dealing with class-imbalanced distributions and spatial consistency constraints. Experiments on the ablation of the two hyperparameters mentioned above will be shown in the Section 4.5. In addition, the number of iterations of the proposed method is set to 100 by default for experimental training.
4.2 Indoor Scene Experiment
Table 3: Quantitative comparison experiment results of semantic segmentation of indoor point cloud dataset ScanNet [45].
@@ 4.2 Indoor Scene Experiment
Table 1 summarises the results for the indoor scene dataset S3DIS [34] Area 5 experimental scenarios. The experimental results are listed as overall and category-level accuracy metrics for the comparison methods [2, 6, 13, 19, 28, 29] and for the proposed method. Bolded data represent the best results for the corresponding metrics, while underlining represents the second best results for the metrics. In the overall indexes, they are better than other comparative methods and state-of-the-art methods, and the proposed method category mIoU are also the best in six of the 13 categories, while the other seven categories have also reached the sub-optimal level. Table 2 summarises the results under the six-fold cross-validation experimental scenario for the indoor scene dataset S3DIS [34]. The overall accuracy, average accuracy, and average cross-validation ratio values of the comparison method [5, 13, 14, 16, 19, 30], as well as the method of this paper, are listed in the experimental results. In addition, the experimental results on the validation set of the ScanNet [45] dataset in Table 3 shows that our method have a great improvement among other methods.
Table 4: Quantitative comparison experiment results of semantic segmentation of outdoor point cloud dataset SemanticKITTI [32] on the validation set. $\ddagger$} represents the results of testing using publicly available pre-trained models.
Method mIoU OA $\ddagger$}
SphereFormer[43] 67.8 91.5
WaffleIron[39] 68.0 -
2DPASS[40] 69.3 -
PTv2[27] 70.3 91.9
OA-CNNs[35] 70.6 92.2
PPT+SparseUNet[50] 71.4 91.6
PTv3+PPT[50, 44] 72.3 92.5
Ours 72.9 92.6
Table 5: Quantitative comparison experiment results of semantic segmentation of outdoor point cloud dataset nuScenes [47].
Method mIoU
OA-CNNs[35] 78.9
FRNet[41] 79.0
WaffleIron[39] 79.1
SphereFormer[43] 79.5
PTv2[27] 80.2
PPT+SparseUNet[50] 78.6
PTv3[50, 44] 80.4
Ours 80.6
Table 6: Quantitative comparison experiment results of semantic segmentation of outdoor point cloud dataset Waymo [48].
Method mIoU mAcc
MinkUNet [36] 65.9 76.6
SPVNAS [37] 67.4 -
Cylinder3D [38] 66.0 -
SphereFormer[43] 69.9 -
PTv2[27] 70.6 80.2
PTv3[44] 71.3 80.5
Ours 71.7 80.9
4.3 Outdoor Scene Experiment
[…]